The total of three sister's ages is 39. Dina is half as old as Michelle and 3 years younger than Juliette. How old are the sisters?
Answer by stanbon(75887)

The total of three sister's ages is 39. Dina is half as old as Michelle and 3 years younger than Juliette. How old are the sisters?
-------
Equations:
d + m + j = 39
d = (1/2)m
d = j-3
---------------
Substitute for m and for j and solve for "d"::
d + 2d + d+3 = 39
4d = 36
d = 9 yrs (Dina's age)
----
m = 2d = 18 yrs (Michelle's age)
----
j = d+3 = 12 yrs (Juliette's age)
-----------
Cheers,
Stan H.
----------